# Morpheus: Benchmarking Physical Reasoning of Video Generative Models with Real Physical Experiments

### v1

@@ Appendix E Prompts for Video Generation Models
Projectile A single, small 3D-printed ball, dark gray in color, is launched from a plastic, small-scale ramp with a slight upward angle. The ball follows a natural, smooth, arcing trajectory upward and then downward, continuing along that arc until it exits the right side of the video frame. The video should accurately simulate the ball’s motion under standard Earth gravity, showing a clear parabolic arc. The ball should not bounce or collide with any objects in the scene. Fixed camera view, no camera movement. In a meticulously crafted scene, a solitary, dark gray 3D-printed ball, with its sleek, spherical form, is propelled from a plastic ramp that slopes gently upward. The ball, weighing a mere fraction of a kilogram, is captured in high-definition, showcasing every nuance of its motion. As it leaves the ramp’s edge, the ball arcs gracefully into the air, its trajectory a perfect parabola that mirrors the laws of physics under standard earth gravity. The video’s frame follows the ball’s smooth ascent and descent, highlighting the ball’s consistent speed and the absence of any sudden accelerations or decelerations. The scene remains unobstructed, ensuring that the ball’s journey is uninterrupted by any external forces, save for the pull of gravity, resulting in a visually stunning and scientifically accurate demonstration of a parabolic motion. Fixed camera view, no camera movement.
Holonomic Pendulum A single pendulum moving retrogressive back and forth. At the bottom of the pendulum, there is a ball attached to it. The pendulum is holonomic. Fixed camera view, no camera movement. A pendulum with a spherical ball attached swings back and forth in a controlled manner, its motion captured in a moment of retrograde swing. The pendulum’s arm, likely made of metal, extends horizontally from a stand, connected to a pivot point that allows for rotational movement. The ball, positioned at the lower end of the pendulum, appears to be in motion, indicating the pendulum’s swing. The environment suggests a laboratory or testing setting, with a backdrop of technical apparatus and equipment, and the lighting is artificial, casting a uniform glow over the scene. The pendulum’s movement, while currently in a retrogressive swing, could potentially change direction, continuing its oscillatory motion. Fixed camera view, no camera movement.
Non-Holonomic Pendulum A single pendulum swings smoothly back and forth. The pendulum consists of a thin, dark string, and at the bottom of the string, there is a small orange ball. The motion of the pendulum is realistic, with slowing at the peaks of its arc and accelerating through the center, simulating gravity’s effect. The pendulum is non-holonomic, so its swing is not perfectly planar, and there might be small, natural deviations in its path. Fixed camera view, no camera movement. In high-definition clarity, a solitary pendulum gracefully arcs through the air, its thin, dark silk string coiling and uncoiling with each oscillation. At the string’s terminus, a small, vibrant orange ball swings with a life-like fluidity, its trajectory punctuated by the subtle slowing at the zenith of its arc and the swift acceleration as it crosses the midpoint. The pendulum’s non-holonomic nature is evident, as it sways slightly off-axis, revealing the gentle, imperceptible wobble that mimics the real-world influence of gravity. Fixed camera view, no camera movement.
Table A4: Base and enhanced textual prompts used for video generation experiments. Enhanced prompts are generated using ChatGLM [31]
and incorporate more detailed scene descriptions and contextual elements derived from the reference images.
Appendix F Additional Analysis
@@
In Figure LABEL:fig:falling_ball_analysis we present additional visualization of the energy and acceleration conversation. In Figure A9 and Figure 9 we visualize difference between real-world and generated videos object trajectories.
F.1 Falling ball
In Figure LABEL:fig:falling_ball_analysis (a), we present the total, kinetic, and potential energy over time. As expected, the total energy dissipates with every new bounce while remaining nearly constant between bounces. Per the video generation method, the discard rate ranges from $11\%$ for COSMOS with enhanced text prompts up to $77\%$ for CogVideoX in a single-frame conditioning regime with plain prompt. The best Dynamical score is achieved by the COSMOS model of $0.35$, which is, however, still far from the real-world video score ($0.99$). Physical invariance scores range from $0.149$ (CogVideoX, single-frame, plain text prompt) to $0.582$ (PyramidalFlow, multi-frame, plain text prompt). We notice that enhanced text prompts increase the scores for COSMOS and CogVideoX, but reduce it for LTX and PyramidalFlow. The qualitative analysis of the discard reasons in Fig. A12 (upper) provides the clue for the low scores, i.e. the overall abundance of duplicates in the videos, as well as the lack of motion. The qualitative inspection of falling and bouncing ball (Fig. LABEL:fig:falling_ball_analysis, c-d) supports this claim. In the case of CogVideoX, generated videos partially exhibit the stillness problem, characterized by an initial absence of motion followed by abrupt and chaotic movements, which deviate significantly from realistic physical behavior.
F.2 Bouncing ball
The main results for the bouncing ball are presented in Tab. A6. Real-world videos achieve optimal performance, reflected by a minimal Discard Rate ($0.0$), a high Dynamical score ($0.99$), and a high Physical Invariance score ($0.99$). In contrast, generated videos display notably lower physical invariance scores, especially evident in plain prompts. The discard rates, presented in Fig.A12 (lower) are lower on average compared to other experiments. Among single-frame methods, COSMOS (enhanced) achieves the highest physical invariance score ($0.546$) but still falls substantially below the real-world baseline. Multi-frame methods generally exhibit improved dynamical and physical scores compared to their single-frame counterparts, although their overall scores remain significantly below those of real-world videos. The first and last frame conditioning with plain prompts surprisingly shows slightly better performance (Physical Invariance score of $0.587$) compared to the enhanced version ($0.550$).
F.3 Projectile
In the case of the projectile, we present the main results in Tab. A7. As for generated videos, the Discard Rates seem to have the same range as for the falling ball, with the main reasons for discarding being the presence of duplicate objects and stillness (see Fig. A12 (middle)), though the disappearance now exhibiting more often, e.g. LTX multi-frame with enhanced prompting. Though the Dynamical and Physical Invariance scores seem $10\%$ lower for the real-world videos compared to other experiments, they are still unreachable for other models.
F.4 Holonomic pendulum
We present the main results in Tab. A8. Real-world videos achieve near-perfect Dynamical ($0.99$) and Physical Invariance ($0.94$) scores, confirming our expectation that a real holonomic pendulum demonstrates consistent, physically accurate periodic motion. The discard rates vary considerably, from as low as $0\%$ for first&frame CogVideoX (enhanced prompts) to as high as $92\%$ for single-frame LTX with enhanced prompts (mainly due to the lack of motion, see Fig. A13 (upper)), indicating severe reliability issues for certain models. Multi-frame configurations tend to outperform single-frame models, achieving higher Dynamical and Physical Invariance scores, with COSMOS (multi-frame, enhanced) and PyramidalFlow (multi-frame, enhanced) showing strong relative performance. Additionally, CogVideoX using the first&last frame conditioning achieves notably high Dynamical scores ($0.85$), though still below the real-world benchmark. Despite capturing some periodic characteristics, the generated methods remain significantly below real-world realism, both quantitatively and qualitatively, suggesting substantial limitations in their ability to accurately model simple pendulum dynamics.
F.5 Non-holonomic pendulum
We present the main results for the non-holonomic pendulum experiment in Tab. A9. The real-world videos achieve near-ideal performance, indicated by a minimal discard rate ($0.000$), a high Dynamical score ($0.98$), and an excellent Physical Invariance score ($0.996$), validating that even non-holonomic pendulum motion adheres closely to expected physical invariants. Generated videos exhibit variable performance. Discard rates are generally lower compared to the holonomic pendulum scenario, suggesting that non-holonomic pendulum dynamics might be somewhat easier for generative models to approximate. CogVideoX, especially in multi-frame setups, demonstrates low discard rates (as low as $0.03$) and high Dynamical scores (up to $0.86$), indicating it can better capture some periodic characteristics of the pendulum. Nevertheless, Physical Invariance scores remain considerably below real-world performance, with the highest being $0.687$ for single-frame COSMOS with plain prompts. Qualitative inspections (Fig. A9) further confirm that while periodic motion patterns, responsible for high Dynamical scores, are partially reproduced, the generated trajectories still significantly deviate from realistic physical behaviors in terms of energy conservation.
F.6 Double pendulum
We present the main results for the double pendulum experiment in Tab. A10. Real-world videos achieve near-perfect Dynamical ($0.99$) and high Physical Invariance scores ($0.938$), serving as a robust benchmark and validating our expectations of minimal dynamical error and strong adherence to physical invariance. Generated videos achieve relatively high Dynamical scores, with the best synthetic performance being $0.94$ (single-frame PyramidalFlow plain prompt). The discard rates vary significantly, from an ideal $0.0\%$ for COSMOS (single-frame, plain) to as high as $77.8\%$ for LTX (single-frame, enhanced), highlighting considerable variability among methods. Despite relatively strong Dynamical scores (up to $0.94$), even the best-performing models show non-negligible dynamical errors (NMSE of $0.06$ compared to the real-world NMSE of $0.002$). The high Dynamical score but low Physical Invariance score may indicate plausible periodic-like motion, but no adherence to the actual physics, see Fig. 9.
F.7 Plain vs. enhanced text prompting
Enhanced text prompting consistently leads to a marked reduction in the discard rate across nearly all experiments and models (e.g., Fig. 6), indicating that the relevance of the generated videos strongly depends on model prompting. However, the gains in terms of the other scores are mixed. We can see that the difference varies from model to model and experiment to experiment, either increasing or decreasing the score or does not provide any significant changes.
Figure A9: Real (left) vs. generated (right) for projectile (upper) and non-holonomic pendulum (lower).
Figure A10: Real (top left) and generated trajectories for the holonomic pendulum.
@@ F.6 Double pendulum
Figure A11: Real (left) and generated trajectory (right) for the double pendulum and corresponding fitting curve with PINN. While NMSE for generated trajectory is small $0.05$, it is still $50$ times worse that PINN with the same parameters fitted to real-world trajectory.
Figure A12: Discard rate reasons for Falling Ball, Projectile and Bouncing Ball experiments
Figure A13: Discard rate reasons for Pendulum experiments
Experiment Name Categories Prompt Type Discard Rate($\downarrow$) Dynamical Score($\uparrow$) Physical Invariance Score($\uparrow$) Energy Conservation($\uparrow$) Acceleration Conservation($\uparrow$) Horizontal Momentum Conservation($\uparrow$)
real-world real-world video - 0.000 0.99 0.977 0.993 0.946 0.994
COSMOS single-frame enhanced 0.109 0.35 0.547 0.617 0.281 0.742
COSMOS single-frame plain 0.347 0.25 0.369 0.417 0.204 0.485
CogVideo single-frame enhanced 0.441 0.12 0.370 0.387 0.245 0.477
CogVideo single-frame plain 0.765 0.05 0.149 0.149 0.102 0.195
LTX single-frame enhanced 0.520 0.09 0.392 0.423 0.295 0.459
LTX single-frame plain 0.280 0.29 0.447 0.470 0.314 0.556
PyramidalFlow single-frame enhanced 0.320 0.26 0.491 0.587 0.299 0.586
PyramidalFlow single-frame plain 0.112 0.18 0.569 0.669 0.358 0.678
COSMOS multi-frame enhanced 0.416 0.20 0.395 0.447 0.202 0.535
COSMOS multi-frame plain 0.673 0.14 0.233 0.269 0.131 0.299
CogVideo multi-frame enhanced 0.400 0.21 0.457 0.463 0.337 0.572
CogVideo multi-frame plain 0.562 0.12 0.320 0.344 0.226 0.390
LTX multi-frame enhanced 0.450 0.18 0.371 0.389 0.243 0.480
LTX multi-frame plain 0.230 0.26 0.444 0.478 0.281 0.573
PyramidalFlow multi-frame enhanced 0.270 0.22 0.511 0.567 0.307 0.659
PyramidalFlow multi-frame plain 0.140 0.34 0.582 0.629 0.381 0.736
CogVideo first&last frame enhanced 0.500 0.13 0.372 0.371 0.308 0.437
CogVideo first&last frame plain 0.714 0.07 0.220 0.230 0.175 0.255
Table A5: Discard Rate, Dynamical score, Physical Invariance score, and Conservation metrics for Falling Ball
Experiment Name Categories Prompt Type Discard Rate($\downarrow$) Dynamical Score($\uparrow$) Physical Invariance Score($\uparrow$) Energy Conservation($\uparrow$) Acceleration Conservation($\uparrow$) Horizontal Momentum Conservation($\uparrow$)
real-world real-world video - 0.000 0.99 0.985 0.991 0.975 0.990
COSMOS single-frame enhanced 0.250 0.05 0.546 0.672 0.272 0.693
COSMOS single-frame plain 0.500 0.17 0.276 0.321 0.162 0.346
CogVideo single-frame enhanced 0.211 0.05 0.457 0.546 0.222 0.605
CogVideo single-frame plain 0.053 0.12 0.509 0.554 0.228 0.744
LTX single-frame enhanced 0.350 0.06 0.333 0.367 0.198 0.434
LTX single-frame plain 0.150 0.07 0.463 0.534 0.211 0.642
PyramidalFlow single-frame enhanced 0.100 0.11 0.663 0.789 0.382 0.819
PyramidalFlow single-frame plain 0.150 0.13 0.602 0.648 0.375 0.782
COSMOS multi-frame enhanced 0.400 0.35 0.362 0.341 0.229 0.516
COSMOS multi-frame plain 0.350 0.35 0.368 0.345 0.229 0.529
CogVideo multi-frame enhanced 0.500 0.11 0.331 0.335 0.237 0.422
CogVideo multi-frame plain 0.250 0.00 0.387 0.397 0.163 0.600
LTX multi-frame enhanced 0.050 0.12 0.476 0.532 0.265 0.630
LTX multi-frame plain 0.150 0.04 0.370 0.405 0.150 0.554
PyramidalFlow multi-frame enhanced 0.150 0.17 0.622 0.641 0.468 0.758
PyramidalFlow multi-frame plain 0.158 0.13 0.620 0.663 0.439 0.757
CogVideo first&last frame enhanced 0.200 0.22 0.550 0.537 0.498 0.614
CogVideo first&last frame plain 0.150 0.26 0.587 0.567 0.461 0.733
Table A6: Discard Rate, Dynamical score, Physical Invariance score, and Conservation metrics for Bouncing Ball
Experiment Name Categories Prompt Type Discard Rate($\downarrow$) Dynamical Score($\uparrow$) Physical Invariance Score($\uparrow$) Energy Conservation($\uparrow$) Acceleration Conservation($\uparrow$) Horizontal Momentum Conservation($\uparrow$)
real-world real-world video - 0.000 0.83 0.862 0.928 0.705 0.952
COSMOS single-frame enhanced 0.574 0.31 0.276 0.327 0.139 0.362
COSMOS single-frame plain 0.435 0.31 0.308 0.382 0.148 0.393
CogVideo single-frame enhanced 0.588 0.31 0.251 0.319 0.164 0.269
CogVideo single-frame plain 0.740 0.18 0.134 0.180 0.087 0.135
LTX single-frame enhanced 0.540 0.41 0.401 0.458 0.291 0.454
LTX single-frame plain 0.420 0.52 0.500 0.571 0.371 0.559
PyramidalFlow single-frame enhanced 0.480 0.43 0.410 0.470 0.282 0.479
PyramidalFlow single-frame plain 0.430 0.44 0.410 0.479 0.275 0.476
COSMOS multi-frame enhanced 0.386 0.30 0.350 0.382 0.209 0.460
COSMOS multi-frame plain 0.402 0.28 0.297 0.305 0.178 0.407
CogVideo multi-frame enhanced 0.430 0.38 0.273 0.354 0.199 0.267
CogVideo multi-frame plain 0.550 0.28 0.227 0.280 0.180 0.220
LTX multi-frame enhanced 0.760 0.19 0.150 0.165 0.118 0.168
LTX multi-frame plain 0.690 0.22 0.168 0.190 0.129 0.186
PyramidalFlow multi-frame enhanced 0.660 0.17 0.232 0.280 0.140 0.277
PyramidalFlow multi-frame plain 0.394 0.34 0.400 0.459 0.255 0.484
CogVideo first&last frame enhanced 0.200 0.61 0.506 0.542 0.335 0.640
CogVideo first&last frame plain 0.530 0.35 0.275 0.301 0.188 0.334
Table A7: Discard Rate, Dynamical score, Physical Invariance score, and Conservation metrics for Projectile
Experiment Name Categories Prompt Type Discard Rate($\downarrow$) Dynamical Score($\uparrow$) Physical Invariance Score($\uparrow$) Energy Conservation($\uparrow$) Period Conservation($\uparrow$) Distance Conservation($\uparrow$)
real-world real-world video - 0.000 0.98 0.939 0.999 0.882 0.936
COSMOS single-frame enhanced 0.370 0.51 0.484 0.477 0.443 0.531
COSMOS single-frame plain 0.250 0.54 0.481 0.607 0.531 0.305
CogVideo single-frame enhanced 0.141 0.35 0.545 0.539 0.331 0.765
CogVideo single-frame plain 0.400 0.43 0.352 0.279 0.312 0.464
LTX single-frame enhanced 0.920 0.05 0.065 0.072 0.047 0.076
LTX single-frame plain 0.560 0.26 0.348 0.407 0.223 0.414
PyramidalFlow single-frame enhanced 0.340 0.43 0.523 0.405 0.579 0.585
PyramidalFlow single-frame plain 0.235 0.45 0.408 0.617 0.315 0.290
COSMOS multi-frame enhanced 0.170 0.67 0.703 0.614 0.691 0.802
COSMOS multi-frame plain 0.374 0.47 0.535 0.475 0.524 0.604
CogVideo multi-frame enhanced 0.070 0.37 0.585 0.368 0.577 0.811
CogVideo multi-frame plain 0.140 0.53 0.528 0.388 0.508 0.688
LTX multi-frame enhanced 0.140 0.50 0.619 0.584 0.509 0.763
LTX multi-frame plain 0.240 0.52 0.599 0.537 0.595 0.663
PyramidalFlow multi-frame enhanced 0.130 0.46 0.686 0.555 0.718 0.785
PyramidalFlow multi-frame plain 0.130 0.45 0.548 0.702 0.517 0.424
CogVideo first&last frame enhanced 0.000 0.85 0.634 0.553 0.555 0.793
CogVideo first&last frame plain 0.030 0.85 0.660 0.613 0.604 0.763
Table A8: Discard Rate, Dynamical score, Physical Invariance score, and Conservation metrics for Holonomic Pendulum
Experiment Name Categories Prompt Type Discard Rate($\downarrow$) Dynamical Score($\uparrow$) Physical Invariance Score($\uparrow$)
real-world real-world video - 0.000 0.98 0.996
COSMOS single-frame enhanced 0.378 0.45 0.537
COSMOS single-frame plain 0.245 0.54 0.687
CogVideo single-frame enhanced 0.101 0.83 0.452
CogVideo single-frame plain 0.061 0.86 0.485
LTX single-frame enhanced 0.720 0.15 0.233
LTX single-frame plain 0.390 0.40 0.504
PyramidalFlow single-frame enhanced 0.330 0.52 0.512
PyramidalFlow single-frame plain 0.300 0.51 0.480
COSMOS multi-frame enhanced 0.364 0.45 0.594
COSMOS multi-frame plain 0.356 0.46 0.609
CogVideo multi-frame enhanced 0.040 0.80 0.503
CogVideo multi-frame plain 0.030 0.80 0.495
LTX multi-frame enhanced 0.130 0.62 0.685
LTX multi-frame plain 0.140 0.58 0.652
PyramidalFlow multi-frame enhanced 0.500 0.34 0.446
PyramidalFlow multi-frame plain 0.480 0.33 0.466
CogVideo first&last frame enhanced 0.230 0.71 0.448
CogVideo first&last frame plain 0.160 0.77 0.503
Table A9: Discard Rate, Dynamical score and Physical Invariance score for Non-nolonomic Pendulum
Experiment Name Categories Prompt Type Discard Rate($\downarrow$) Dynamical Score($\uparrow$) Physical Invariance Score($\uparrow$)
real-world real-world video - 0.000 0.99 0.938
COSMOS single-frame enhanced 0.103 0.86 0.555
COSMOS single-frame plain 0.086 0.86 0.538
CogVideo single-frame enhanced 0.047 0.90 0.578
CogVideo single-frame plain 0.000 0.93 0.537
LTX single-frame enhanced 0.778 0.21 0.152
LTX single-frame plain 0.337 0.64 0.448
PyramidalFlow single-frame enhanced 0.207 0.77 0.463
PyramidalFlow single-frame plain 0.154 0.81 0.511
COSMOS multi-frame enhanced 0.101 0.88 0.547
COSMOS multi-frame plain 0.134 0.83 0.528
CogVideo multi-frame enhanced 0.060 0.88 0.526
CogVideo multi-frame plain 0.000 0.91 0.566
LTX multi-frame enhanced 0.059 0.91 0.617
LTX multi-frame plain 0.015 0.94 0.604
PyramidalFlow multi-frame enhanced 0.149 0.84 0.565
PyramidalFlow multi-frame plain 0.152 0.83 0.551
Table A10: Discard Rate, Dynamical score and Physical Invariance score for Double Pendulum
Appendix G Impact statement
Advancing physics-aware generative models is crucial for bridging perception and reasoning in AI, enabling more reliable simulations for robotics, scientific discovery, and autonomous systems. This paper introduces MORPHEUS, a benchmark revealing significant physical violations in video generation models, highlighting the need for integrating physics constraints into generative AI.

### v2

@@ Appendix F Video Generative Models Details
Collision Generate a realistic video of two metallic spheres colliding. In the first frames the leftmost sphere in the video is moving towards the static one. At the moment of contact the ball in motion transfer its kinetic energy to the ball at rest. The two spheres have identical physical properties, namely material, shapes, masses. The momentum should be conserved. Fixed camera view, no camera movement. In a high-definition, slow-motion sequence, two identical metallic spheres, each polished to a mirror-like finish, are meticulously positioned on a frictionless surface. The left sphere, propelled by an unseen force, hurtles towards the stationary sphere, its trajectory a perfect parabola. As the oncoming sphere approaches, the static sphere begins to subtly vibrate, a prelude to the impending collision. The moment of contact is captured in stunning detail, revealing the transfer of kinetic energy as the moving sphere’s momentum is transferred to the stationary one. The spheres, crafted from the same dense material, maintain their spherical shapes and masses, ensuring the conservation of momentum throughout the impact. The scene is illuminated by a single, soft light source, casting long shadows and emphasizing the physics of the collision. Fixed camera view, no camera movement.
Spring A single metallic slotted mass attached on a spring moving periodically up and down. Fixed camera view, no camera movement. In a meticulously crafted scene, a solitary metallic slotted mass, polished to a gleaming silver, is ingeniously attached to a tensioned spring. The mass, weighing several pounds, oscillates with a fluid grace, its movement initiated by the subtle release of the spring. The camera captures the mass as it arcs upwards, the tension in the spring visible, before it gently descends with a rhythmic sway, the sound of its metallic slats clinking softly in the background. The scene is set against a stark white backdrop, emphasizing the stark contrast between the mass and the spring, and the smooth, periodic motion of the mechanical dance. Fixed camera view, no camera movement.
Table 6: Base and enhanced textual prompts used for video generation experiments: Double Pendulum, Bouncing Ball, Collision, Spring. Enhanced prompts are generated using corresponding upsampler VLMs such as ChatGLM (Zeng et al., 2024) and incorporate more detailed scene descriptions and contextual elements derived from the reference images. Slight modifications to these prompts were made in case of different variants such as object type.
Appendix G Prompts for Video Generation Models
For each experiment, we carefully designed a prompt that describes the physical setup and motion of the experiment being conducted. For example, in the falling ball experiment, the prompt specifies that the ball falls and makes contact with the table below. Similarly, in the projectile experiment, we describe how the ball is launched at a slight upward angle and follows a natural parabolic trajectory. We enhance these prompts using an internally provided upsampler or ChatGLM if no upsampler is provided to incorporate more detailed scene descriptions and contextual elements derived from the reference images. All prompts are shown in Table 5 and Table 6.
Appendix H Additional Analysis
Figure 16: The distribution of the average physical invariance score across models and experiments.
Figure 17: The distribution of the average dynamical score across models and experiments.
Figure 18: The distribution of the average total score across models and experiments.
Figure 19: The distribution of the average discard rate due to the objects’ disappearance across models and experiments.
Figure 20: The distribution of the average discard rate due to the objects’ duplicates across models and experiments.
Figure 21: The distribution of the average discard rate due to objects’ stillness across models and experiments.
Figure 22: The relative change in scores evaluating on original data vs. on the augmented data. The augmented examples seem to be much harder.
Figure 23: The relative change in scores for using an enhanced prompt vs the plain one.
Figure 24: The scores for the holonomic pendulum experiment. On the left, darker colors denote enhanced textual prompt.
@@ Appendix H Additional Analysis
In Figure 26(a), we present an additional visualization of the energy and acceleration conservation. In Figure 15, we visualize the difference between the object trajectories of real-world and generated videos. In Figure 26(a)(a), we present the total, kinetic, and potential energy over time. As expected, the total energy dissipates with every new bounce while remaining nearly constant between bounces.
(a) Energy analysis of real-world and generated falling + bouncing ball videos: (a) Real-world video energy conservation (b) CogVideoX plain single frame generated video energy conservation
Appendix I Usage of Large Language Models
Large language models were used solely for grammar correction and language polishing of the manuscript. They had no involvement in the research design, experimentation, analysis, or generation of results; all technical contributions are the work of the authors.
Ethics statement.
This work involves the collection of benchmark datasets and the evaluation of publicly available models, all used in compliance with their original licenses. No new human data or personal annotations were gathered. Our study focuses exclusively on benchmarking and reproducibility, without any collection or processing of user data.
Reproducability statement.
For reproducibility, we release the full code on GitHub and all benchmark data on Hugging Face.
Figure 26: Real (top left) and generated trajectories for the holonomic pendulum.
@@ Appendix I Usage of Large Language Models
Figure 27: Real (left) and generated trajectory (right) for the double pendulum and corresponding fitting curve with PINN. While NMSE for generated trajectory is small $0.05$, it is still $50$ times worse that PINN with the same parameters fitted to real-world trajectory.